(no subject)

arenton@uvahee.phys.virginia.edu
Sun, 3 Oct 1999 08:04:07 -0400

Shift Summary for owl shift, 03/Oct/99

On Shift: Mike Arenton, Ed Blucher

Hours of beam delivered at all intensities: 6.2/7
Hours of beam delivered at nominal intensity: 6.2
Hours of E799 data (special runs not included): 5.8
Hours of special runs ( muon runs...):0.1

Accounting of lost time: tape change, magnet reversals, run starts

========================================================================
Runs ended this shift:

Run 14795 - Beam 799 - 25 M events
Run 14796 - Beam muon
Run 14797 - Beam 799 - ongoing

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift:none

========================================================================
Overview of shift:
E799 data taking, did muon run and flipped NM4AN and NM2SR
polarity to NEG.

While changing tapes discovered 5 UPC tapes ready to be used
were actually old DLT III (10 Gbyte) tapes (brown cases) rather
than DLT IIIXT (15 Gbyte) (white cases) that we are using now.
These were replaced with XT's and a check shows no other old
tapes in the labled store.

Spill structure is quite variable and generally worse than
24 hours ago.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14795 ??? 00:10 CSI 2603 & 2610 have PED=0, pipeline reset
with DBC fixes it. Don't know spill number
because DAMP display is frozen.

00:30 - 00:40 no beam

14795 385 00:45 per Sasha L's request changed TRD HV's to
values expected to put MIP peak at 35 ADC cnts.
set usrlatch to 1.

14795 02:20 DAMP, including the mechanism that determines
when the tape is full, appears not to be
updating. Notice DOT warnings of tape full,
manually stop run. Calibration job stops,
but rest of DA continues to run - stop triggers
save histo's, do fresh start with new tapes.

14796 02:30 start muon run, flip NM4AN, NM2SR polarity to NEG.

14797 02:50 start BEAM_799 run, most mon_disp displays not
working, restart them with fix_disp.

03:00 notice that TRD buffer gas system is shut down,
has been this way since 17:00 yesterday.
Sasha G is called.

14797 03:30 - 03:40 beam off, RF problem, when beam returns
spill structure, which has been poor, is
much better, this only lasts for about
30 minutes, however.

03:40 Sasha G. has got TRD buffer gas system
running again.

04:00 onsplit mount UPK007 on ktev1

06:05 onsplit mount UPF014 on ktev4

06:45 beam off, no explanation
07:00 beam returns

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift